OpenStudy (anonymous):

Help with L'Hopital's Rule The problem I'm working on: lim (cotx)^x x->0+ I would really like a step by step if possible, I'm completely stumped on this problem!

zepdrix (zepdrix):

\(\Large\bf \color{#008353}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

\[\Large\bf\sf \lim_{x\to0^+}(\cot x)^x\]So if you try to evaluate this directly, it looks like our function is approaching the indeterminate form: \(\Large\bf\sf \infty^0\)

zepdrix (zepdrix):

Soooo we'll have to try something fancy.

OpenStudy (anonymous):

That's where we swap e with the lim, right?

zepdrix (zepdrix):

No, because the base is `approaching` a large number. The exponent is `approaching` zero. But which one is getting there faster? We can't necessary say that it's approaching a number to the zero power.

zepdrix (zepdrix):

Yah plugging in a small small number is a useful :) Usually works. But let's do the L'Hop method as the problem asks. Yes, we have to do a sneaky little trick involving e. \[\Large\bf\sf e^{\ln x}\quad=\quad x\]Does that look familiar?

OpenStudy (anonymous):

Yes, so far I have: \[e^\lim xlncotx \] as x approaches 0+

zepdrix (zepdrix):

Ok good. Let's ignore the e for right now, and just pay attention to what's going on up there.\[\Large\bf\sf \lim_{x\to0^+} x \ln(\cot x)\]

zepdrix (zepdrix):

We have to do another weird little trick from here :)

OpenStudy (anonymous):

that's an I.F. lim as x->0+ ln(cotx)/(1/x)

zepdrix (zepdrix):

i.f?

OpenStudy (anonymous):

Indeterminate Form

zepdrix (zepdrix):

True. But it's now the I.F that we want, right? With the way we have it written now, the function is approaching \(\Large\bf\sf \dfrac{\infty}{\infty}\)

zepdrix (zepdrix):

So we can apply L'Hop to it from this form!

OpenStudy (anonymous):

Okay, first L'Hop! (csc^2x/cotx)/(-1/x^2)

OpenStudy (anonymous):

This is where I start getting really lost

zepdrix (zepdrix):

Negative sign on the top somewhere also yes? (cot x)' = -csc^2x I think.

OpenStudy (anonymous):

oops! You're right!

zepdrix (zepdrix):

Hmm yah things are getting tricky here aren't they? We can convert everything to sines and cosines. It might help.

zepdrix (zepdrix):

I'm ending up with this limit now,\[\Large\bf\sf \lim_{x\to0^+}\frac{x^2}{\sin x \cos x}\]I went through it quickly though, possible error. So try it and see if you get the same thing.

OpenStudy (anonymous):

Let's see. . .1/cotx = tanx, or sinx/cosx and -csc^2x = 1/sin^2x, so we have (x^2)\sinxcosx YES!

OpenStudy (anonymous):

Oops, I forgot a negative in the 1\sin^2x

zepdrix (zepdrix):

yah it cancel's with the negative from x^2. No big deal.

zepdrix (zepdrix):

Rewrite the denominator using Sine Double Angle Identity (will make it easier to differentiate).

zepdrix (zepdrix):

\[\Large\bf\sf \lim_{x\to0^+}\frac{x^2}{\frac{1}{2}\sin(2x)}\]

zepdrix (zepdrix):

And from here we can see that our function is approaching the indeterminate form 0/0. Ok good! Still a good form that we can work with. Time for L'Hop #2!!!

zepdrix (zepdrix):

Trying to figure out where that 1/2 came from? XD lol

OpenStudy (anonymous):

Yes! I'm terrible with the double angles and half angles!

zepdrix (zepdrix):

Here is our Double Angle Identity,\[\Large\bf\sf \color{royalblue}{\sin(2x)=2\sin x \cos x}\]If we divide both sides by 2,\[\Large\bf\sf \frac{1}{2}\sin(2x)=\sin x \cos x\]The sinx cosx is the part we're replacing in our integral.

OpenStudy (anonymous):

Okay, I see it now!

OpenStudy (anonymous):

Let's see if I can differentiate. . .

OpenStudy (anonymous):

2x/[1/2 cos(2x) 2] => 2x\cos(2x)

zepdrix (zepdrix):

Ok good.

zepdrix (zepdrix):

Are we still getting an indeterminate form as x->0 ?

OpenStudy (anonymous):

No!, now we have 0/1

zepdrix (zepdrix):

Ooo interesting! So our exponent ( the whole thing we had up inside of the e ) = 0.

zepdrix (zepdrix):

So our initial guess was correct! :) lol

OpenStudy (anonymous):

AHA! e^0 = 1!

OpenStudy (anonymous):

Thanks so much! You really are awesome!

OpenStudy (anonymous):

:)

zepdrix (zepdrix):

Np \c:/